This deals with perimeters and area from a practice worksheet, not a textbook
I am not sure how to solve:
1) A right triangle with base 5 inches, height 12 inches and hypotenuse 13 inches.
2)A circle with a diameter 14 inches
Thank you for your help. If there is a formula for these equations, what is it?

Answer by stanbon(48535)

This deals with perimeters and area from a practice worksheet, not a textbook
I am not sure how to solve:
1) A right triangle with base 5 inches, height 12 inches and hypotenuse 13 inches.
Area = (1/2)base*height
= (1/2)5*12 = 30 sq. in.
Perimeter + 5+12+13 = 30 in
-----------------------------------------
2)A circle with a diameter 14 inches
Area = pi*radius^2
A + pi*7^2
Area = 49*pi sq. in.
Perimeter = pi*diameter
P = 14pi inches
========================
Cheers,
Stan H.